Prev Next

SW&IT Meeting Minutes 2021-09-13

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20210906
LR20210906 - Bec, Matthieu 40.0
LR20210906 - Mechtley, Brandon 0.0
LR20210906 - Eychaner, Glenn 32.0
LR20210906 - Edmeades, David 0.0
LR20210906 - Grenz, Paul 19.0
LR20210906 - Hooper, Stephen 0.0
LR20210906 - Zhang, Yang 32.0
LR20210906 - Zhang, Xianyu 0.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

  • dms CA's expiring tomorrow

Team reports

MB

  • Last week
    • OC, Management, ATT PM
    • AO s/w KoM meet-up Aracteri
    • IT8470 - guider jumps, detector problem? will continue today
    • IT8462 - horizon queries, timeout from poor network
    • IT7882 - FACSUM wfs images, fixed mount on info.mountain
    • AO SX investigations - masterdiag CBs, filotto stats 2018-2021, open some new github issues, follow-up With Brandon (moving files)
    • archive meeting, agree on inotify scheme, discussed follow-up with Yang
  • This week
    • SWG reports, overdue
    • zip ATT delivered folder, add new docs to working locations

GCE
  • Last week
    • Work around git filename case issues on macOS with a case-sensitive disk image.
    • Continued study of IIF and DMS/Telemetry.
    • Work on integration of ahs-luci into the global ahs system.
    • Review of open LUCI IssueTrak issues. (There are 99 at last count; 19 are marked as High and 11 as Software, but some may be misclassified).
    • Continue monitoring IT 8454 (disk space on luci.luci).
    • Finish benefits enrollment.
  • This week
    • Hope the mountain network issue has been repaired. It makes working on mountain computers difficult.
    • Test (with Jenny Mon. and Tues. night), replicate, and begin to fix the highest-priority LUCI issues (the scripting and RTD in particular).
    • Finish integrating the ahs-luci code into the global ahs system.

DE

  • Windows server, no response from CDW despite pings. I may just get the install media and try a test install to see if there is a tool included.
  • Backup server parts on their way
  • ME offboarding, defer email account until 11 Oct
  • ME laptop wipe and check in. As discussed, there were frequent performance problems which may or may not be solved by the wipe and new OS. I have marked it so that we can deploy it appropriately.
  • Archive meeting/Cristina says she will handle the new year directory, but I am watching it in case she doesn't have time to put a real solution in place
  • AC wired network issue: I reregistered his machine as lab equipment from my admin account, have not heard of a recurrence of the problem.
  • JW managed to break his monitor setup, am attempting to understand what he did.
  • Vacation rest of the week.

YZ

  • Previous:
    • LBC bias-level adjustments with Alessandro and Olga (for newly installed controller #4, the Red Sci CCD).
    • modsAlign problem fixes and some oddities (don't yet understand why) reported by Olga.
    • "lbc_cmu" problem fixes on robs with MB and SH.
    • DMS refmem refactoring, new WCS client, and some discussion with MB. (PR merged.)
    • Splitting MODS github repo started (mods-obs repo created).
    • TMS meeting and some possible new software work proposed by Andrew (regarding M1 thermo).
    • Some discussion of new image notification utility to be run on archive machine with MB.
  • Next:
    • Some misc MODS/LBC problems to investigate (modsAlign for example).
    • Start to work on MODS github repo reorganization if have time.
    • Discuss with Olga a recent MODS IDL problem that OSU hasn't responded yet.
    • Investigate new image notification utility with Python 2.7.
    • TMS software work planning, and perhaps read Andrew's new proposal.

BM

  • Last week:
    • Some work parsing logs for actuator currents before and after overcurrent events.
    • More reading re: AdSec hardware design/software interface.
    • Meetings / activities:
      • AOG overcurrent meeting.
      • AO Q&As with Al: More overcurrent history, briefly discussed remote observer client.
      • Arcetri re: AO software handover, power backplane, SWG learning plan.
      • Ride along with Juan Carlos and Gustavo for ARGOS tests (cut short due to network problems.)
  • This week:
    • A few MasterDiagnostic logging issues in Github.
    • Prevent adaptive secondary .bin files from being deleted before being backed up and migrate AO logs from Mountain to Tucson.
    • Meetings / activities:
      • AOG meeting re: Gain=0 testing (by Friday?)
      • JCG/GR ARGOS testing Friday.
      • AO sessions with Al.
    • Research / learning:
      • Continue overcurrent (circular buffer), power backplane upgrade research.
      • WFS magnitude calculation in AOS code

PSG

  • Last week:
    • Finally tested the ARGOS swing arm behavior from tcs1.
    • Connectivity is ok.
    • Commands to the swing arms worked, as did responses. Swing arms moved correctly.
  • This week:
    • Testing and debugging the final version of the code in the OSS.

SH

  • vacation last Tuesday
  • Mountain network issues
  • PEPSI KVM call
  • base camp camera
  • mountain DNS (infrastructure machines)

next:

  • mountain LDAP, RADIUS
  • Mountain network issues
Topic revision: r12 - 13 Sep 2021, GlennEychaner
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback